# H-AMR FORGE’d in FIRE I: Magnetic state transitions, jet launching and radiative emission in super-Eddington, highly magnetized quasar disks formed from cosmological initial conditions

### v2

@@ 1 Introduction
In a recent series of papers (Hopkins et al., 2024a, b, c, henceforth “H24”, collectively), the authors simulated the formation of a quasar accretion disk from first principles and incorporated a variety of multiphysics (e.g., radiation magnetohydrodynamics, thermochemistry, star formation). H24 accomplished this by beginning with cosmological initial conditions, zooming in on an individual galaxy, and then using the Lagrangian hyper-refinement method to resolve the nascent quasar disk down to $\gtrsim 300\,r_{\rm g}$, where $r_{\rm g}\equiv GM/c^{2}$ is the gravitational radius of the SMBH. The resulting disk was distinct from tradiational accretion disks: it cooled rapidly on sub-orbital timescales and was dominated by toroidal magnetic fields with plasma $\beta\equiv p_{\rm t}/p_{\rm b}\ll 1$, where $p_{\rm t}$ and $p_{\rm B}$ are the gas thermal and magnetic pressures, respectively. These properties result from the mechanism of disk assembly: the SMBH accretes dynamically cold, already-magnetized gas from the interstellar medium (ISM), resulting in streams of gas that circularize with preferentially strong, $p_{\rm B}\gg p_{\rm t}$, toroidal magnetic fields supplied by the advected toroidal magnetic flux, which, in turn, drives rapid, super-Eddington accretion. This is also quite similar to Gaburov et al. (2012), who performed the first of such simulations of SMBH disks formed from the disruption of cold, magnetized interstellar gas, and found similarly dominant toroidal magnetic flux. In both H24 and Gaburov et al. (2012), these strong magnetic fields stabilized the gas from star formation (as specifically focused on in Hopkins et al., 2024a), resolving the long-standing issue of gravitational fragmentation in classical accretion disk models.
Recently, Hopkins et al. (2024d) developed models of magnetically dominated disks (MDDs) to explain this new type of accretion flow analytically (see also Johansen & Levin, 2008). These MDDs are distinct from magnetically arrested disks (MADs, Bisnovatyi-Kogan & Ruzmaikin, 1974; Igumenshchev et al., 2003; Narayan et al., 2003; Tchekhovskoy et al., 2011), which are defined by having strong net vertical magnetic flux, but still have strong toroidal fields (although low net toroidal flux) and maintain $\beta\sim 1$ in their midplanes. Despite the nomenclature “arrested”, MADs actually accrete extremely efficiently through a combination of magnetized turbulence and winds (e.g., Manikantan et al., 2024), leading to inflow within a tenth of an orbital timescale (Igumenshchev et al., 2003; Narayan et al., 2003; Jacquemin-Ide et al., 2021; Scepi et al., 2024). Traditional MADs are radiatively inefficient and have usually been studied without explicit cooling or radiation. However, the MAD state may also extend to more luminous regimes where the net vertical magnetic flux is saturated (e.g., Avara et al., 2016; Curd & Narayan, 2023; Scepi et al., 2024; Liska et al., 2024, Lowell et al. in prep 2024).
@@ 2.2 Divergence Cleaning
Once we have interpolated all quantities onto the H-AMR grid, they have been faithfully translated from GIZMO up to the truncation error of the interpolation scheme. This truncation error is acceptably small for all quantities except for the magnetic fields. This is because the divergence of the magnetic fields, $\vec{\nabla}\cdot\vec{B}$, must be kept as close to zero as possible to reliably perform magnetohydrodynamic calculations. H-AMR uses a staggered mesh to evolve magnetic fields and conserves the initial value of $\vec{\nabla}\cdot\vec{B}$ to machine precision. After the interpolation, the maximum value of $|\vec{\nabla}\cdot\vec{B}|$ is $\approx 6000$ in code units (see Section 2.3) and $\approx 0.1$ in normalized units ($\Delta x|\vec{\nabla}\cdot\vec{B}/B|$, where $\Delta x$ is the average cell width) . This pollutes our solution with unphysical magnetic monopoles, which we must “clean” to obtain trustworthy results. Divergence cleaning usually involves an elliptic equation, which must be solved over the entire grid simultaneously. This is prohibitively expensive and especially complicated on our spherical, non-uniform grids with static and adaptive mesh refinement. To make progress, we have introduced a novel version of “scalar” divergence cleaning (e.g., Balsara & Kim, 2004), which splits the global problem up into many small local problems (similar in spirit, but different in practice, is a cell-by-cell method introducd by Silberman et al., 2019). We describe this approach in detail in Appendix B, where we have verified that the maximum value of $\Delta x|\vec{\nabla}\cdot\vec{B}/B|$ after cleaning is $\mathcal{O}(10^{-7})$. Importantly, our cleaning approach largely leaves the interpolated magnetic field the same. We provide a map of $|\vec{\nabla}\cdot\vec{B}|$ before and after cleaning in Figure 16 in the Appendix.
@@ 2.3 H-AMR Simulation Setup
Code, grid and boundary conditions. H-AMR (Liska et al., 2022) is a GPU-accelerated, three-dimensional, general-relativistic, radiation magnetohydrodynamics (GRRMHD) code. We use a static metric and a spherical polar grid that is centered on the BH, where the base grid is uniform in the modified Kerr-Schild coordinates, ${\rm log}r$, $\theta$ and $\varphi$. The inner grid radius is sufficiently deep inside the event horizon ($r_{\rm in}=0.9\,r_{\rm H}$) such that the region beyond the event horizon is causally disconnected from the inner radial grid boundary. The outer radius of the simulation is at $r_{\rm out}\approx 10^{5}\,r_{\rm g}$ where
@@
Radiation treatment. H-AMR uses the two-moment “M1” closure scheme for radiation (Levermore, 1984), which posits that there is always a frame in which the radiation field is isotropic. We separately evolve the radiation energy density, $E_{\rm r}$ (in the radiation frame), and the radiation velocity, $u_{\rm r}^{\mu}$ (which is the first moment of the frequency-integrated intensity). We operator split the Riemann problem to separately calculate the radiative and non-radiative fluxes. We evolve radiation using a second-order implicit-explicit (IMEX) time integrator, wherein the radiative flux is calculated during the explicit part of the timestep. We use absorption-averaged opacities that are valid from $T\gtrsim 10^{4}\,{\rm K}$ to $T\lesssim 10^{10}\,{\rm K}$ and the high densities (relative to, e.g., the interstellar medium) modeled in the disk. The opacities include free-free, bound-free and bound-bound processes fit over a range of density and temperatures, a Gaussian profile in temperature at $\approx 1.5\times 10^{5}\,{\rm K}$ to treat the iron line opacity, synchrotron opacities fitted for SMBHs, and electron scattering opacities which account for thermal Comptonization, as described in McKinney et al. (2014, 2017). Our opacities assume local thermodynamic equilibrium and depend on $\rho$, gas temperature $T_{\rm g}$, and radiation temperature $T_{\rm r}$. We assume thermal radiation such that $T_{\rm r}=(\hat{E}_{\rm r}/a)^{1/4}$, where $\hat{E}_{\rm r}$ is the fluid-frame radiation energy density and $a$ is the radiation constant. The gas temperature is given by $T_{\rm g}=(\gamma-1)u_{\rm g}\mu_{\rm g}m_{\rm H}c^{2}/k_{\rm B}\rho$, where we use the mean molecular weight $\mu_{\rm g}=\frac{4}{6X+Y+2}\approx 0.61$ (where $X=0.70$ and $Y=0.28$ are the adopted hydrogen and helium mass fractions, respectively) and we set $\gamma=5/3$ whenever radiation is activated. While or opacities are reliable within the optically thick body of the disk, they are not well-motivated beyond the photosphere (which, we will see, is very large) or at the cold ($\lesssim 10^{4}\,{\rm K}$) outer edges of our disk. However, the outer disk does not dynamically evolve in the runtime of our simulations, so they remain consistent with the GIZMO simulation that treated the thermodynamics of these regions reliably. Our photosphere does evolve, but the thermodynamics at the photosphere does not affect the dynamics of the accretion flow itself.
@@
Sim. Name BH Spin Non-radiative Runtime Radiative Runtime
@@
Simulations. In this work, we report two simulations, which differ in the value of the dimensionless BH spin, $a$. We label the simulations “non-spinning” (NS), where $a=0$, and “highly-spinning” (HS), where $a=0.9375$; see Table 1. We adopt a base grid of $1536$ logarithmically-spaced radial cells, $384$ polar cells, and $128$ azimuthal cells. Using 1D SMR, we gradually increase the azimuthal resolution to $512$ azimuthal cells within 60 degrees of the equator. We also use “internal derefinement”, which gradually decreases the azimuthal resolution within the grid blocks adjacent to the poles (for details see Liska et al., 2022). This results in roughly cubical cells that resolve the characteristic density scale height of the disk, $(H/r)_{\rho}\sim 0.2-0.5$, by $\sim 24-60$ cells.
@@ 3.3 Radial profiles
Thermal, radiation and magnetic pressures. In Fig. 5(b), we show the ratios of the different pressures in the disk. Each component of the pressure is density-weighted to focus the average on the disk region (i.e., we show ratios of $\langle p_{i}\rangle_{\rm disk}$, where $i$ indicates the pressure component). The pressure ratios in HS and NS are nearly identical. The thermal pressure, $p_{\rm t}$, is highly subdominant to both the magnetic pressure, $p_{\rm B}$, and the radiation pressure, $p_{\rm r}$, everywhere. However, $p_{\rm B}$ and $p_{\rm r}$ are more comparable, and in the inner disk ($\lesssim 200\,r_{\rm g}$), $p_{\rm B}\approx p_{\rm r}$. In the outer disk, $p_{\rm B}$ exceeds $p_{\rm r}$ by a factor $\approx 10-10^{2}$. This is roughly consistent with H24, who generally find $p_{\rm B}\approx 10^{2}p_{\rm r}$ at large radii. This is also consistent with the analytic predictions of Hopkins et al. (2024d) and Hopkins (2024).
@@ 3.4.2 What causes the magnetic state transition?
Once the poloidal field becomes structured and dipolar, it quickly transforms toroidal field in kind. This is because the orbital shear generates a toroidal field of odd parity from the dipolar poloidal field on an orbital timescale. In the outer disk, the even parity, quadrupole-like toroidal field is supplied mainly by flux-freezing (see arguments in Hopkins et al., 2024c), which occurs on an inflow timescale (the Parker dynamo likely also plays a role, e.g. Johansen & Levin, 2008; Gaburov et al., 2012; Squire et al., 2024). Therefore, the parity of the toroidal field will be determined by a competition between the advection of quadrupole-like $\mathcal{B}_{\varphi}^{\rm S}$ and the shear of dipole-like $\mathcal{B}_{r}^{\rm S}$ and $\mathcal{B}_{\theta}^{\rm S}$. The ratio of these two generation mechanisms (shear $v_{\phi}$ vs. advection $v_{r}$) is thus a condition for the dominant parity. Indeed, if the condition,
@@
A similar NVF region also develops in Jacquemin-Ide et al. (2024), where in a GRMHD simulation of a torus with an initially toroidal field, the advection of intermediate-scale turbulent poloidal field loops enabled the development of a strong, dipolar, poloidal field in the inner $\gtrsim 300\,r_{\rm g}$. In this work, the toroidal field evolution depended strongly on magnetic buoyancy. Magnetic buoyancy may eject toroidal field lines and thus reduce the strength of the NTF in the outer disk. However, NTF may also sustained when $\beta<1$ if a Parker dynamo regenerates the field (Johansen & Levin, 2008; Squire et al., 2024). In H24, the toroidal field was sustained by continued inwards advection of NTF from the tidally disrupted giant molecular cloud that formed the disk (Hopkins et al., 2024c). Magnetic buoyancy presumably plays an important role in our disk as well, but requires a more detailed analysis that is beyond the scope of this paper.
@@ 3.5.1 The structure of the outflow and emission regions
The white contours in Figure 10 show the axisymmetrized effective photosphere ($\tau_{\rm eff}=1$) averaged over both hemispheres, which extends to $r\gtrsim 10^{4}\,r_{\rm g}$. In HS, the BZ jet carves out a cavity, allowing the effective photosphere to reach down to the BH. This is important, as it suggests that emission from the ISCO – including, for instance, the iron K-$\alpha$ line which is often used to infer the BH spin (Reynolds, 2021) – can escape at very high viewing angles. If viewed face-on, this system would likely appear blazar-like. However, at more intermediate viewing angles, the observed emission is significantly reprocessed by the time it escapes into the optically thin regions. In NS, there is no BZ jet, and the effective photosphere is at $r\gtrsim 10^{4}\,r_{\rm g}$ even directly above the BH. This indicates that all emission escaping the disk will be reprocessed. We also note the flared structure of the photosphere at $x\approx 5-15\times 10^{3}\,r_{\rm g}$, which likely plays an important role in shaping the broad-line region (e.g., Hopkins, 2024).

### v4

@@ 1 Introduction
In a recent series of papers (Hopkins et al., 2024a, c, d, henceforth “H24”, collectively), the authors simulated the formation of a quasar accretion disk from first principles and incorporated a variety of multiphysics (e.g., radiation magnetohydrodynamics, thermochemistry, star formation). H24 accomplished this by beginning with cosmological initial conditions, zooming in on an individual galaxy, and then using the Lagrangian hyper-refinement method to resolve the nascent quasar disk down to $\gtrsim 300\,r_{\rm g}$, where $r_{\rm g}\equiv GM/c^{2}$ is the gravitational radius of the SMBH. The resulting disk was distinct from tradiational accretion disks: it cooled rapidly on sub-orbital timescales and was dominated by toroidal magnetic fields with plasma $\beta\equiv p_{\rm t}/p_{\rm b}\ll 1$, where $p_{\rm t}$ and $p_{\rm B}$ are the gas thermal and magnetic pressures, respectively. These properties result from the mechanism of disk assembly: the SMBH accretes dynamically cold, already-magnetized gas from the interstellar medium (ISM), resulting in streams of gas that circularize with preferentially strong, $p_{\rm B}\gg p_{\rm t}$, toroidal magnetic fields supplied by the advected toroidal magnetic flux, which, in turn, drives rapid, super-Eddington accretion. This is also quite similar to Gaburov et al. (2012), who performed the first of such simulations of SMBH disks formed from the disruption of cold, magnetized interstellar gas, and found similarly dominant toroidal magnetic flux. In both H24 and Gaburov et al. (2012), these strong magnetic fields stabilized the gas from star formation (as specifically focused on in Hopkins et al., 2024a), resolving the long-standing issue of gravitational fragmentation in classical accretion disk models.
Recently, Hopkins et al. (2024e) developed models of magnetically dominated disks (MDDs) to explain this new type of accretion flow analytically (see also Johansen & Levin, 2008). These MDDs are distinct from magnetically arrested disks (MADs, Bisnovatyi-Kogan & Ruzmaikin, 1974; Igumenshchev et al., 2003; Narayan et al., 2003; Tchekhovskoy et al., 2011), which are defined by having strong net vertical magnetic flux, but still have strong toroidal fields (although low net toroidal flux) and maintain $\beta\sim 1$ in their midplanes. Despite the nomenclature “arrested”, MADs actually accrete extremely efficiently through a combination of magnetized turbulence and winds (e.g., Manikantan et al., 2024), leading to inflow within a tenth of an orbital timescale (Igumenshchev et al., 2003; Narayan et al., 2003; Jacquemin-Ide et al., 2021; Scepi et al., 2024). Traditional MADs are radiatively inefficient and have usually been studied without explicit cooling or radiation. However, the MAD state may also extend to more luminous regimes where the net vertical magnetic flux is saturated (e.g., Avara et al., 2016; Curd & Narayan, 2023; Scepi et al., 2024; Liska et al., 2024, Lowell et al. in prep 2024).
@@ 2.2 Divergence Cleaning
Once we have interpolated all quantities onto the H-AMR grid, they have been faithfully translated from GIZMO up to the truncation error of the interpolation scheme. This truncation error is acceptably small for all quantities except for the magnetic fields. This is because the divergence of the magnetic fields, $\vec{\nabla}\cdot\vec{B}$, must be kept as close to zero as possible to reliably perform magnetohydrodynamic calculations. H-AMR uses a staggered mesh to evolve magnetic fields and conserves the initial value of $\vec{\nabla}\cdot\vec{B}$ to machine precision. After the interpolation, the maximum value of $|\vec{\nabla}\cdot\vec{B}|$ is $\approx 6000$ in code units (see Section 2.3) and $\approx 0.1$ in normalized units ($\Delta x|\vec{\nabla}\cdot\vec{B}/B|$, where $\Delta x$ is the average cell width) . This pollutes our solution with unphysical magnetic monopoles, which we must “clean” to obtain trustworthy results. Divergence cleaning usually involves an elliptic equation, which must be solved over the entire grid simultaneously. This is prohibitively expensive and especially complicated on our spherical, non-uniform grids with static and adaptive mesh refinement. To make progress, we have introduced a novel version of “scalar” divergence cleaning (e.g., Balsara & Kim, 2004), which splits the global problem up into many small local problems (similar in spirit, but different in practice, is a cell-by-cell method introducd by Silberman et al., 2019). We describe this approach in detail in Appendix B, where we have verified that the maximum value of $\Delta x|\vec{\nabla}\cdot\vec{B}/B|$ after cleaning is $\mathcal{O}(10^{-7})$. Importantly, our cleaning approach largely leaves the interpolated magnetic field the same. We provide a map of $|\vec{\nabla}\cdot\vec{B}|$ before and after cleaning in Figure 16 in the Appendix. The combined process of interpolation and divergence cleaning takes a few hours for our problem.
@@ 2.3 H-AMR Simulation Setup
Code, grid and boundary conditions. H-AMR (Liska et al., 2022) is a GPU-accelerated, three-dimensional, general-relativistic, radiation magnetohydrodynamics (GRRMHD) code. We use a stationary metric and a spherical polar grid that is centered on the BH, where the base grid is uniform in the modified Kerr-Schild coordinates, ${\rm log}r$, $\theta$ and $\varphi$. The inner grid radius is sufficiently deep inside the event horizon ($r_{\rm in}=0.9\,r_{\rm H}$) such that the region beyond the event horizon is causally disconnected from the inner radial grid boundary. The outer radius of the simulation is at $r_{\rm out}\approx 10^{5}\,r_{\rm g}$ where
@@
Radiation treatment. H-AMR uses the two-moment “M1” closure scheme for radiation (Levermore, 1984), which posits that there is always a frame in which the radiation field is isotropic. We separately evolve the radiation energy density, $E_{\rm r}$ (in the radiation frame), and the radiation velocity, $u_{\rm r}^{\mu}$ (which is the first moment of the frequency-integrated intensity). We operator split the Riemann problem to separately calculate the radiative and non-radiative fluxes. We evolve radiation using a second-order implicit-explicit (IMEX) time integrator, wherein the radiative flux is calculated during the explicit part of the timestep. We use absorption-averaged opacities that are valid from $T\gtrsim 10^{4}\,{\rm K}$ to $T\lesssim 10^{10}\,{\rm K}$ and the high densities (relative to, e.g., the interstellar medium) modeled in the disk. The opacities include free-free, bound-free and bound-bound processes fit over a range of density and temperatures, a Gaussian profile in temperature at $\approx 1.5\times 10^{5}\,{\rm K}$ to treat the iron line opacity, synchrotron opacities fitted for SMBHs, and electron scattering opacities which account for thermal Comptonization, as described in McKinney et al. (2014, 2017). Our opacities assume local thermodynamic equilibrium and depend on $\rho$, gas temperature $T_{\rm g}$, and radiation temperature $T_{\rm r}$. We assume thermal radiation such that $T_{\rm r}=(\hat{E}_{\rm r}/a)^{1/4}$, where $\hat{E}_{\rm r}$ is the fluid-frame radiation energy density and $a$ is the radiation constant. The gas temperature is given by $T_{\rm g}=(\gamma-1)u_{\rm g}\mu_{\rm g}m_{\rm H}c^{2}/k_{\rm B}\rho$, where we use the mean molecular weight $\mu_{\rm g}=\frac{4}{6X+Y+2}\approx 0.61$ (where $X=0.70$ and $Y=0.28$ are the adopted hydrogen and helium mass fractions, respectively) and we set $\gamma=5/3$ whenever radiation is activated. While our opacities are reliable within the optically thick body of the disk, they are not well-motivated beyond the photosphere (which, we will see, is very large) or at the cold ($\lesssim 10^{4}\,{\rm K}$) outer edges of our disk. However, the outer disk does not dynamically evolve in the runtime of our simulations, so they remain consistent with the GIZMO simulation that treated the thermodynamics of these regions reliably. Our photosphere does evolve, but the thermodynamics at the photosphere does not affect the dynamics of the accretion flow itself.
@@
Sim. Name BH Spin Non-Rad. Runtime Rad. Runtime
@@
Simulations. In this work, we report two simulations, which differ in the value of the dimensionless BH spin, $a$. We label the simulations “non-spinning” (NS), where $a=0$, and “highly-spinning” (HS), where $a=0.9375$; see Table 1. We adopt a base grid of $1536$ logarithmically-spaced radial cells, $384$ polar cells, and $128$ azimuthal cells. Using 1D SMR, we gradually increase the azimuthal resolution to $512$ azimuthal cells within 60 degrees of the equator. We also use “internal derefinement”, which gradually decreases the azimuthal resolution within the grid blocks adjacent to the poles (for details see Liska et al., 2022). This results in roughly cubical cells that resolve the characteristic density scale height of the disk, $(H/r)_{\rho}\sim 0.2-0.5$, by $\sim 24-60$ cells. Usually, magnetized disk simulations are considered resolved if their cell size is sufficiently small compared to the wavelength of the magnetorotational instability (Hawley et al., 2011). Since our disk is extremely magnetized, this wavelength is on the order of the disk scale height itself, so we consider our resolution adequate.
@@ 3.3 Radial profiles
Thermal, radiation and magnetic pressures. In Fig. 5(b), we show the ratios of the different pressures in the disk. Each component of the pressure is density-weighted to focus the average on the disk region (i.e., we show ratios of $\langle p_{i}\rangle_{\rm disk}$, where $i$ indicates the pressure component). The pressure ratios in HS and NS are nearly identical. The thermal pressure, $p_{\rm t}$, is highly subdominant to both the magnetic pressure, $p_{\rm B}$, and the radiation pressure, $p_{\rm r}$, everywhere. However, $p_{\rm B}$ and $p_{\rm r}$ are more comparable, and in the inner disk ($\lesssim 200\,r_{\rm g}$), $p_{\rm B}\approx p_{\rm r}$. In the outer disk, $p_{\rm B}$ exceeds $p_{\rm r}$ by a factor $\approx 10-10^{2}$. This is roughly consistent with H24, who generally find $p_{\rm B}\approx 10^{2}p_{\rm r}$ at large radii. This is also consistent with the analytic predictions of Hopkins et al. (2024e) and Hopkins et al. (2024b).
@@ 3.4.2 What causes the magnetic state transition?
Once the poloidal field becomes structured and dipolar, it quickly transforms toroidal field in kind. This is because the orbital shear generates a toroidal field of odd parity from the dipolar poloidal field on an orbital timescale. In the outer disk, the even parity, quadrupole-like toroidal field is supplied mainly by flux-freezing (see arguments in Hopkins et al., 2024d), which occurs on an inflow timescale (the Parker dynamo likely also plays a role, e.g. Johansen & Levin, 2008; Gaburov et al., 2012; Squire et al., 2024). Therefore, the parity of the toroidal field will be determined by a competition between the advection of quadrupole-like $\mathcal{B}_{\varphi}^{\rm S}$ and the shear of dipole-like $\mathcal{B}_{r}^{\rm S}$ and $\mathcal{B}_{\theta}^{\rm S}$. The ratio of these two generation mechanisms (shear $v_{\phi}$ vs. advection $v_{r}$) is thus a condition for the dominant parity. Indeed, if the condition,
@@
A similar NVF region also develops in Jacquemin-Ide et al. (2024), where in a GRMHD simulation of a torus with an initially toroidal field, the advection of intermediate-scale turbulent poloidal field loops enabled the development of a strong, dipolar, poloidal field in the inner $\gtrsim 300\,r_{\rm g}$. In this work, the toroidal field evolution depended strongly on magnetic buoyancy. Magnetic buoyancy may eject toroidal field lines and thus reduce the strength of the NTF in the outer disk. However, NTF may also sustained when $\beta<1$ if a Parker dynamo regenerates the field (Johansen & Levin, 2008; Squire et al., 2024). In H24, the toroidal field was sustained by continued inwards advection of NTF from the tidally disrupted giant molecular cloud that formed the disk (Hopkins et al., 2024d). Magnetic buoyancy presumably plays an important role in our disk as well, but requires a more detailed analysis that is beyond the scope of this paper.
@@ 3.5.1 The structure of the outflow and emission regions
The white contours in Figure 10 show the axisymmetrized effective photosphere ($\tau_{\rm eff}=1$) averaged over both hemispheres, which extends to $r\gtrsim 10^{4}\,r_{\rm g}$. In HS, the BZ jet carves out a cavity, allowing the effective photosphere to reach down to the BH. This is important, as it suggests that emission from the ISCO – including, for instance, the iron K-$\alpha$ line which is often used to infer the BH spin (Reynolds, 2021) – can escape at very high viewing angles. If viewed face-on, this system would likely appear blazar-like. However, at more intermediate viewing angles, the observed emission is significantly reprocessed by the time it escapes into the optically thin regions. In NS, there is no BZ jet, and the effective photosphere is at $r\gtrsim 10^{4}\,r_{\rm g}$ even directly above the BH. This indicates that all emission escaping the disk will be reprocessed. We also note the flared structure of the photosphere at $x\approx 5-15\times 10^{3}\,r_{\rm g}$, which likely plays an important role in shaping the broad-line region (e.g., Hopkins et al., 2024b).